OpenStudy (anonymous):

A barrel contains 189 gallons of water and is being drained at a constant rate of 5 gallons per hour. Write an equation that models the number of gallons g, after t hours. Can anyone help me with this equation, or an example of one similar?

OpenStudy (mathstudent55):

At the beginning, when t = 0, g = 189

OpenStudy (mathstudent55):

As time goes by, when t = 1, it lost 5 gallons times 1, or -5 gallons t = 2, it lost 5 gallons times 2, or -5 * 2 = -10 gallons t = t, it lost 5 gallons times t, or -5t gallons The expression is the original amount plus the loss: g = 189 - 5t

OpenStudy (mathstudent55):

To check to see if this expression is correct, substtitute a few values of t and evaluate it. At t = 0, no water has drained, so it should have the full 189 gallons: g = 189 - 5(0) = 189 - 0 = 189 After 1 hour, at t = 1, it lost 5 gallons, so it has 189 - 5 = 184: g = 189 - 5(1) = 189 - 5 = 184 It looks like g = 189- 5t is the correct expression.

OpenStudy (anonymous):

I see... That makes sense! Thank you very much. I wasn't sure how to do this, but that makes a lot of sense now. Thank you, thank you!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

one sec-

OpenStudy (anonymous):

when you were figuring it out, after you wrote t=1, t=2,etc., you wrote t=t, it lost 5 gallons times t or -5 gallons..

OpenStudy (anonymous):

is that just the expression w/out plugging in an amount for t?

OpenStudy (mathstudent55):

t is the general term for t number of hours. That's how we get the -5t part of the expression. In 1 hour it looses 5 gal In 2 hours it looses 10 gal In 3 hours it looses 15 gal In t hours it looses 5t gal

OpenStudy (anonymous):

oh, ok. Duh me!! LOL thanks!